Circular No. 7875

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2002ce IN NGC 2604 R. Arbour, South Wonston, Hants., England, reports his discovery

of an apparent supernova ∼ 16′′ west and 12′′ south of the nucleus of the barred-spiral galaxy NGC 2604. The object, of mag 16.0–16.5, was found on an unfiltered CCD image obtained with a 0.3-m f/6.3 Schmidt-Cassegrain and Starlight Xpress CCD camera on Apr. 10.878 UT and confirmed on Apr. 11.863. The object’s existence was also confirmed on Apr. 10.909 and 11.838 by T. Boles, Coddenham, Suffolk, who gave the position as α = 8^{h}33^{m}22^{s}.1, δ = +29^{o}32′01′′ (equinox 2000.0). Arbour notes that the object was not present on his image of the region on Apr. 5.200, and it is also absent on the second-generation Digitised Sky Survey.

T. Matheson, S. Jha, P. Challis and R. Kirshner, Harvard-Smithsonian Center for Astrophysics, report that a spectrum (range 370–750 nm), obtained by P. Berlind on Apr. 12.18 UT with the F. L. Whipple Observatory 1.5-m telescope (+ FAST spectrograph), shows it to be a type II supernova near maximum. The spectrum consists of a blue continuum with P-Cyg lines of hydrogen. Adopting the NASA/IPAC Extragalactic Database recession velocity of 2094 km/s for the host galaxy, the expansion velocity derived from the minimum of the Hβ line is 5300 km/s.

SUPERNOVA 2002ca IN UGC 8521 Matheson et al. also report that a similar spectrum, obtained by Ber-

lind on Apr. 5.35 UT, also having a blue continuum and P-Cyg lines of hydrogen, shows SN 2002ca (cf. IAUC 7866) to be a type II supernova before maximum. Adopting recession velocity 3277 km/s for the host galaxy, the expansion velocity derived from the minimum of the Hβ line is 7900 km/s.

COMETS C/2002 A4 AND C/2002 G3 (SOHO) Further to IAUC 7873, D. Hammer and D. Biesecker report their mea-

surements of comets found by R. Kracht on C2 and by X.-M. Zhou on C3 coronagraph images on the SOHO website. Reduced astrometry and orbital elements by the undersigned appear on the MPEC s cited below. C/2002 A4 belongs to the Meyer group. C/2002 G3, estimated at m_{1} = 8.5 ± 0.5 and perhaps at r = 0.27 AU at discovery, should brighten some in SOHO images as it approaches perihelion (r = q = 0.08 AU) on Apr. 17.

Comet  2002 UT  α_{2000}  δ_{2000}  MPEC
C/2002 A4 Jan. 1.163  18^{h}48.^{m}8  −21^{o}56′  2002-G45
C/2002 G3 Apr. 12.113  1 40.2  + 1 54  2002-G51
2002 April 12  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Brian G. Marsden